# Characterization of the Cherenkov Photon Background for Low-Noise Silicon Detectors in Space

### v1

@@ 5 RESULTS
Note. — Cherenkov rates assume cosmic-ray flux at L2 during median solar activity and the specified masking radius. The Cherenkov rate for the MAS CCD includes Cherenkov photons generated during the 30 s readout time.
@@
We simulate the cosmic-ray environment at L2 for two detector configurations (Table 5). The first configuration mimics the Teledyne-e2v CCD201-20 EMCCD detectors that were initially baselined for the Roman Coronagraph Instrument (CGI; Spergel et al., 2015; Harding et al., 2016). We note that for these simulations, we care only about detector thickness and pixel size, and thus our results should be applicable to newer EMCCD detector architectures (i.e., the CCD301, CCD302, and CCD311; Morrissey et al., 2023). These are relatively thin detectors (13 m thick) with fast readout and good quantum efficiency in the visible. The detector characteristics including readout noise, clock-induced charge, and dark current were collected by Lacy et al. (2019) based on measurements from Harding et al. (2016). We chose these detectors because they were used for the exoplanet sensitivity study presented in Lacy et al. (2019).
@@
The second detector configuration corresponds to a future red-optimized silicon detector with a thickness of m. This detector configuration is envisioned as a future iteration of the -channel MAS CCD (Holland, 2023; Botti et al., 2024), which inherits the radiation tolerance, low readout noise, and quantum efficiency of the skipper CCD (e.g., Dawson et al., 2008; Tiffenberg et al., 2017; Marrufo Villalpando et al., 2024; Roach et al., 2024; Cervantes-Vergara et al., 2025). The MAS CCD achieves deeply sub-electron readout noise through the repeated measurement of the charge in each pixel by a series of floating gate skipper amplifiers. We envision a detector with , m pixels read through two serial registers, each containing 300 non-destructive amplifiers with an inter-amplifier separation of 3 pixels. Such a device could be read in 30 s in frame-transfer mode with a readout speed of kpix/s (Lapi et al., 2025; Lin et al., 2025). Assuming a single-sample readout noise of 2.5 e- rms/pix (Cancelo et al., 2021), this would result in a final readout noise of e- rms/pix. We converted the readout noise into the rate of readout-noise-induced spurious charge reported in Table 5. We note that this spurious charge rate comes from sampling the exponential tail of the readout noise distribution and is very sensitive to the exact value of the readout noise. This rate could be reduced by improving the single-sample noise, adding more output amplifiers, or measuring the charge multiple times per amplifier (Lapi et al., 2025). Similar readout noise could also be achieved over smaller regions-of-interest in a comparable readout time using fewer amplifiers (Chierchie et al., 2021). Furthermore, a thick, -channel CCD read with multiple SiSeRO amplifiers could achieve similar readout noise and speed with fewer output channels (Sofo-Haro et al., 2024). We take values for clock-induced charge and dark current from Barak et al. (2022); however, this performance has yet to be demonstrated in an astronomy-optimized detector.
@@ 5.1 Detector Thickness
The residual Cherenkov background rate depends on detector thickness. Thinner detectors reduce Cherenkov emission by limiting the amount of material traversed by high-energy particles leading to shorter tracks, fewer Cherenkov photons, and a lower probability that Cherenkov photons are absorbed by the detector. However, this suppression comes at the cost of reduced quantum efficiency at longer wavelengths. This trade-off is critical when targeting spectral features such as the absorption band at 940 nm, which is of particular interest in the search for biosignatures (e.g., Seager, 2013; Rauscher et al., 2019).
In order to isolate the effect of detector thickness, we start from the EMCCD configuration in Table 5 and artificially change its thickness while keeping all other detector parameters fixed. For each detector thickness, we adjust the corresponding quantum efficiency and effective exposure time after masking. Figure 6 shows how the residual Cherenkov background and quantum efficiency change with detector thickness and masking radius. We compute a simple SNR considering literature values for intrinsic detector backgrounds (Table 5) and a nominal signal rate of 0.0016 e-/pix/s from Lacy et al. (2019). We find higher SNR for thicker detectors, indicating that the increased quantum efficiency dominates over the increased residual Cherenkov background. Furthermore, our analysis suggests that a smaller masking radius maximizes SNR because a larger masking radius reduces the total effective exposure time. We note that the upper bound of a 20 pix masking radius is much larger than conventional cosmic-ray masking algorithms (e.g., van Dokkum, 2001) and leads to a significant () reduction in total signal integration time.
@@ 5.2 Exoplanet Sensitivity
We proceed to include our model for the residual Cherenkov photon background in the Roman CGI exoplanet atmosphere sensitivity model described by Lacy et al. (2019). Lacy et al. (2019) provide a detailed description and open source code111https://github.com/blacy/direct-imaging-sims for estimating signal and background sources for exoplanet spectroscopy, with a focus on the giant exoplanet observations with the Roman CGI. The Lacy et al. (2019) model includes intrinsic detector backgrounds (dark current, clock induced charge, and readout noise), as well as speckle, zodiacal, and exozodiacal light.
In Figure 7, we present estimates for the signal and background rates for the Roman CGI operating in integral field spectrograph (IFS) mode for an observation of a Jupiter-radius planet orbiting a 5.0 mag G0V star at a separation of 3.8 AU observed at a distance of 10 pc. Using the Lacy et al. (2019) model for a 13-m-thick Teledyne-e2v CCD201-20 EMCCD, we find that the residual Cherenkov background rate is subdominant to other detector and astronomical backgrounds, as well as the exoplanet signal. However, this thin detector provides minimal sensitivity at longer wavelengths, with at 940 nm after 410 hours of integrated exposure time. In contrast, the envisioned 250-m-thick fully depleted MAS CCD has a Cherenkov background rate that is larger than other intrinsic detector backgrounds and is comparable to the exoplanet signal and astronomical backgrounds. This configuration yields at 940 nm after 410 hours of integrated exposure time. We note that assuming a 30 s readout time per exposure in the MAS CCD configuration doubles the dark current and residual Cherenkov contributions relative to the signal. A 2-pixel masking radius is assumed for the residual Cherenkov background rate presented in Figure 7, while the gray band represents the variation expected from solar modulation. Variations in solar activity affect both the residual Cherenkov background rate and the fraction of the detector that is masked due to cosmic-ray tracks.
@@ 6 CONCLUSIONS
We emphasize that Cherenkov radiation is an irreducible source of background for all silicon detectors in space, as well as any other detectors using dielectric materials where Cherenkov radiation can be emitted and detected (Du et al., 2022).
@@
The cosmic-ray track length, and thus the Cherenkov photon rate, can be reduced with thinner detectors; however, thin detectors are less sensitive at long wavelengths, which is a critical regime for identifying biosignatures in the spectra of exoplanet atmospheres (Rauscher et al., 2022).
@@ Acknowledgments
We thank David Maurin for helpful discussions about the cosmic-ray environment at L2. This work was partially funded by NASA APRA (No. 80NSSC22K1411), Fermilab LDRD (2019.011, 2022.054), and the Heising-Simons Foundation (#2023-4611).
@@ Appendix A Contribution of Heavy Nuclei
As noted before, heavy nuclei make a negligible contribution to the overall cosmic-ray rate. While it is expected that heavy nuclei will also contribute negligibly to the Cherenkov photon rate, the correspondingly larger electric charge of heavy nuclei means that more Cherenkov photons are produced per cosmic-ray event (i.e., the dependence in Eq. 1). From our full simulations, we see that heavy nuclei indeed make a minimal contribution to the residual Cherenkov background rate (Figure 8). This does not mean that heavy nuclei are unimportant to observations; their tracks occupy more pixels than light nuclei, reducing the effective exposure time per pixel after processing. However, this is suppressed for longer exposure times, since protons dominate the detector occupation and overlap between tracks becomes more likely to occur. Thus, we conclude that heavy nuclei can be ignored when estimating the Cherenkov background rate and detector masking; however, the greater charge and more complex event morphologies of heavy nuclei can affect image calibration pipelines (e.g., Regan, 2024).
@@ Appendix B Cherenkov Rates
We note that the rates presented here differ from those calculated in Gaido et al. (2024) due to changes in the cosmic-ray proton rate, how readout time was factored into the observation time, and how the per-frame rate was computed. We believe that the rates presented here represent a more accurate estimate of the residual Cherenkov background rate expected in L2.

### v2

@@ 5 RESULTS
Note. — Cherenkov rates assume cosmic-ray flux at L2 during median solar activity and the specified masking radius. The Cherenkov rate for the MAS CCD includes Cherenkov photons generated during the 30 s readout time. For the EMCCD, we use the parameter values presented in Lacy:2019.
References. — (1) Lacy:2019 (2) Barak:2022, (3) Cancelo:2021, (4) Marrufo:2024.
@@
We simulate the cosmic-ray environment at L2 for two detector configurations (Table 1). The first configuration mimics the Teledyne-e2v CCD201-20 EMCCD detectors that were initially baselined for the Roman Coronagraph Instrument (CGI; Spergel:2015; Harding:2016). We note that for these simulations, we care only about detector thickness and pixel size, and thus our results should be applicable to newer EMCCD detector architectures (i.e., the CCD301, CCD302, and CCD311; Morrissey:2023). These are relatively thin detectors (13 m thick) with fast readout and good quantum efficiency in the visible. The detector characteristics including readout noise, clock-induced charge, and dark current were collected by Lacy:2019 based on measurements from Harding:2016. We chose these detectors because they were used for the exoplanet sensitivity study presented in Lacy:2019.
@@
The second detector configuration corresponds to a future red-optimized silicon detector with a thickness of m. This detector configuration is envisioned as a future iteration of the -channel MAS CCD (Holland:2023; Botti:2024), which inherits the radiation tolerance, low readout noise, and quantum efficiency of the skipper CCD (e.g., Dawson:2008; Tiffenberg:2017; Marrufo:2024; Roach:2024; Cervantes-Vergara:2025). The MAS CCD achieves deeply sub-electron readout noise through the repeated measurement of the charge in each pixel by a series of floating gate skipper amplifiers. We envision a detector with , m pixels read through two serial registers, each containing 300 non-destructive amplifiers with an inter-amplifier separation of 3 pixels. Such a device could be read in 30 s in frame-transfer mode with a readout speed of kpix/s (Lapi:2025; Lin:2025). Assuming a single-sample readout noise of 2.5 e- rms/pix (Cancelo:2021), this would result in a final readout noise of e- rms/pix. We converted the readout noise into the rate of readout-noise-induced spurious charge reported in Table 1. We note that this spurious charge rate comes from sampling the exponential tail of the readout noise distribution and is very sensitive to the exact value of the readout noise. This rate could be reduced by improving the single-sample noise, adding more output amplifiers, or measuring the charge multiple times per amplifier (Lapi:2025). Similar readout noise could also be achieved over smaller regions-of-interest in a comparable readout time using fewer amplifiers (Chierchie:2021). Furthermore, a thick, -channel CCD read with multiple SiSeRO amplifiers could achieve similar readout noise and speed with fewer output channels (Sofo-Haro:2024). We take values for clock-induced charge and dark current from Barak:2022; however, this performance has yet to be demonstrated in an astronomy-optimized detector.
@@ 5.1 Detector Thickness
The residual Cherenkov background rate depends on detector thickness. Thinner detectors reduce Cherenkov emission by limiting the amount of material traversed by high-energy particles leading to shorter tracks, fewer Cherenkov photons, and a lower probability that Cherenkov photons are absorbed by the detector. However, this suppression comes at the cost of reduced quantum efficiency at longer wavelengths. This trade-off is critical when targeting spectral features such as the absorption band at 940 nm, which is of particular interest in the search for biosignatures (e.g., Seager:2013; Rauscher:2019).
In order to isolate the effect of detector thickness, we start from the EMCCD configuration in Table 1 and artificially change its thickness while keeping all other detector parameters fixed. For each detector thickness, we adjust the corresponding quantum efficiency and effective exposure time after masking. Figure 6 shows how the residual Cherenkov background and quantum efficiency change with detector thickness and masking radius. We compute a simple SNR considering literature values for intrinsic detector backgrounds (Table 1) and a nominal signal rate of 0.0016 e-/pix/s from Lacy:2019. We find higher SNR for thicker detectors, indicating that the increased quantum efficiency dominates over the increased residual Cherenkov background. Furthermore, our analysis suggests that a smaller masking radius maximizes SNR because a larger masking radius reduces the total effective exposure time. We note that the upper bound of a 20 pix masking radius is much larger than conventional cosmic-ray masking algorithms (e.g., vanDokkum:2001) and leads to a significant () reduction in total signal integration time.
@@ 5.2 Exoplanet Sensitivity
We proceed to include our model for the residual Cherenkov photon background in the Roman CGI exoplanet atmosphere sensitivity model described by Lacy:2019. Lacy:2019 provide a detailed description and open source code111https://github.com/blacy/direct-imaging-sims for estimating signal and background sources for exoplanet spectroscopy, with a focus on the giant exoplanet observations with the then-current design of the Roman CGI.222Since the analysis of Lacy et al. (2019), the Roman-CGI design moved to a slit-prism spectrograph. This reduces the dark current contribution due to a decrease in the number of pixels covered by the signal. The Lacy:2019 model includes intrinsic detector backgrounds (dark current, clock-induced charge, and readout noise), as well as speckle, zodiacal, and exozodiacal light.
In Figure 7, we present estimates for the signal and background rates for the Roman CGI operating in integral field spectrograph (IFS) mode for an observation of a Jupiter-radius planet orbiting a 5.0 mag G0V star at a separation of 3.8 AU observed at a distance of 10 pc. Using the Lacy:2019 model for a 13-m-thick Teledyne-e2v CCD201-20 EMCCD, we find that the residual Cherenkov background rate is subdominant to other detector and astronomical backgrounds, as well as the exoplanet signal. However, this thin detector provides minimal sensitivity at longer wavelengths, with at 940 nm after 410 hours of integrated exposure time. In contrast, the envisioned 250-m-thick fully depleted MAS CCD has a Cherenkov background rate that is larger than other intrinsic detector backgrounds and is comparable to the exoplanet signal and astronomical backgrounds. This configuration yields at 940 nm after 410 hours of integrated exposure time. We note that assuming a 30 s readout time per exposure in the MAS CCD configuration doubles the dark current and residual Cherenkov contributions relative to the signal. A 2-pixel masking radius is assumed for the residual Cherenkov background rate presented in Figure 7, while the gray band represents the variation expected from solar modulation. Variations in solar activity affect both the residual Cherenkov background rate and the fraction of the detector that is masked due to cosmic-ray tracks.
@@ 6 CONCLUSIONS
We emphasize that Cherenkov radiation is an irreducible source of background for all silicon detectors in space, as well as any other detectors using dielectric materials where Cherenkov radiation can be emitted and detected (Du:2022).
@@
The cosmic-ray track length, and thus the Cherenkov photon rate, can be reduced with thinner detectors; however, thin detectors are less sensitive at long wavelengths, which is a critical regime for identifying biosignatures in the spectra of exoplanet atmospheres (Rauscher:2022).
@@
Software: Geant4 (Geant4:2002), matplotlib (Hunter:2007), numpy (Harris:2020), scipy (Virtanen:2020).
@@ Acknowledgments
We thank David Maurin for helpful discussions about the cosmic-ray environment at L2. We thank Brianna Lacy for a useful discussion about the Roman CGI instrument model. This work was partially funded by NASA APRA (No. 80NSSC22K1411), Fermilab LDRD (2019.011, 2022.054), and the Heising-Simons Foundation (#2023-4611).
@@ Appendix A Contribution of Heavy Nuclei
As noted before, heavy nuclei make a negligible contribution to the overall cosmic-ray rate. While it is expected that heavy nuclei will also contribute negligibly to the Cherenkov photon rate, the correspondingly larger electric charge of heavy nuclei means that more Cherenkov photons are produced per cosmic-ray event (i.e., the dependence in Eq. 1). From our full simulations, we see that heavy nuclei indeed make a minimal contribution to the residual Cherenkov background rate (Figure 8). This does not mean that heavy nuclei are unimportant to observations; their tracks occupy more pixels than light nuclei, reducing the effective exposure time per pixel after processing. However, this is suppressed for longer exposure times, since protons dominate the detector occupation and overlap between tracks becomes more likely to occur. Thus, we conclude that heavy nuclei can be ignored when estimating the Cherenkov background rate and detector masking; however, the greater charge and more complex event morphologies of heavy nuclei can affect image calibration pipelines (e.g., Regan:2024).
@@ Appendix B Cherenkov Rates
We note that the rates presented here differ from those calculated in Gaido:2024 due to changes in the cosmic-ray proton rate, how readout time was factored into the observation time, and how the per-frame rate was computed. We believe that the rates presented here represent a more accurate estimate of the residual Cherenkov background rate expected in L2.